# Differentiation, the Exception Not the Rule - Evidence for Full Miscibility in Sub-Neptune Interiors

### v1

@@ 8 Acknowledgements
E.D.Y. acknowledges financial support from NASA grant 80NSSC24K0544 (Emerging Worlds program).
@@
In order to include Fe in our mixtures, we rely on linear mixing of densities to define $\rho$ such that
$\frac{1}{\rho}=\frac{w_{\rm sil\,mix}}{\rho_{\rm sil\,mix}}+\frac{1-w_{\rm sil\,mix}}{\rho_{\rm Fe\,melt}},$ (14)
where $w_{\rm sil\,mix}$ is the weight fraction of the melt composed of the silicate-H2 mixture, $\rho_{\rm sil\,mix}$ is the density of that mixture, and $\rho_{\rm Fe\,melt}$ is the density of liquid Fe metal at the same conditions.
@@
The envelope is divided into three regions. The bulk of the lower region is convective. Because the vapor and silicate melt are in equilibrium, we use the moist pseudoadiabat of Graham et al. (2021) to evaluate the convective temperature gradient $\nabla T_{\mathrm{conv}}$. Molar heat capacities required to evaluate the pseudoadiabat from Graham et al. (2021) are obtained from the NIST thermodynamic data base (Chase, 1998) where we assume the vapor silicate component speciates to SiO, Mg, and O2. Hindrance of convection due to the mass load of heavy elements at relatively high temperatures must be included, as described previously for similar circumstances by Misener et al. (2023).
By including the Ledoux criterion rather than just the Schwarzschild criterion for convection, the calculations permit development of a radiative layer at the base of the atmosphere that transports the intrinsic heat coming from the magma ocean upward (Guillot, 2010; Misener et al., 2023). The intrinsic luminosity due to the heat emanating from the core to space becomes a determining factor for the structure of the atmosphere. Since radiative diffusion relies on the local luminosity, the temperature gradient due to radiative diffusion includes the radially-dependent luminosity, $L(r)$. At depths well below the radiative-convective boundary, stellar radiation cannot penetrate the atmosphere. Therefore, at these depths the radiative diffusion is given by
@@
We define the radius of the planet using the chord optical depth as formulated by Guillot (2010) coupled with the opacities from Freedman et al. (2014) with metallicities corresponding to the composition of the gas (essentially pure H2 at relevant heights for the outer atmosphere). In general the radii defined this way are slightly larger than those corresponding to 1 bar the two vary by less than 10%).

### v4

@@ 8 Acknowledgements
We acknowledge financial support from NASA grant 80NSSC24K0544 (Emerging Worlds program).
@@
We model Fe dissolved in a supercritical silicate–H2 melt by assigning Fe a partial molar volume that reflects Fe occupation of Mg-like cation sites within the silicate melt framework. We let $\phi\in[0,1]$ denote the fraction of the pure MgSiO3 formula-unit molar volume attributable to the Mg site, and let $\alpha=(r_{\mathrm{Fe}}/r_{\mathrm{Mg}})^{3}$ scale these site volumes by the ratio of octahedrally-coordinated ionic radii (Shannon, 1976). If $\bar{V}_{\mathrm{MgSiO_{3}}}^{\mathrm{pure}}={\rm MW}_{\mathrm{MgSiO_{3}}}/\rho_{\mathrm{MgSiO_{3}}}^{\mathrm{pure}}$ is the molar volume of pure MgSiO3 at the local $P$ and $T$, the Fe-site partial molar volume is
$\bar{V}_{\mathrm{Fe\,site}}^{\mathrm{eff}}=\phi\,\alpha\;\bar{V}_{\mathrm{MgSiO_{3}}}^{\mathrm{pure}}(P,T),$ (15)
and the corresponding effective partial Fe density is
$\rho_{\mathrm{Fe\,site}}^{\mathrm{eff}}=\frac{{\rm MW}_{\mathrm{Fe}}}{\bar{V}_{\mathrm{Fe\,site}}^{\mathrm{eff}}}.$ (16)
Assuming additivity of specific volumes (no excess mixing volume), the density of the miscible melt with Fe mass fraction $w_{\mathrm{Fe}}$ is given by the harmonic mean
$\frac{1}{\rho_{\mathrm{mix}}}=\frac{1-w_{\mathrm{Fe}}}{\rho_{\mathrm{MgSiO_{3}+H_{2}}}}+\frac{w_{\mathrm{Fe}}}{\rho_{\mathrm{Fe}}^{\mathrm{eff}}},$ (17)
where $\rho_{\mathrm{MgSiO_{3}+H_{2}}}$ is the density of the MgSiO3+H2 mixture derived at $P$ and $T$. An empirical value for $\phi$ of 0.5 is obtained from the MgO-Fe system (Insixiengmay & Stixrude, 2025)). We use the MgO-Fe rather than the MgSiO3-Fe simulations because in the former case NPT ensemble simulations were used to derive volumes at constant pressure, affording the required volumetric data for mixing. In the models reported here we use $r_{\mathrm{Fe}}=0.078~\text{nm}$ (VI, high spin), and $r_{\mathrm{Mg}}=0.072~\text{nm}$. Results are insensitive to the precise values for these constants.
@@
The envelope is divided into three regions. The bulk of the lower region is convective. Because the vapor and silicate melt are in equilibrium, we use the moist pseudoadiabat of Graham et al. (2021) to evaluate the convective temperature gradient $\nabla T_{\mathrm{conv}}$. Molar heat capacities required to evaluate the pseudoadiabat from Graham et al. (2021) are obtained from the NIST thermodynamic data base (Chase, 1998) where we assume the vapor silicate component speciates to SiO, Mg, and O2. If we scale the pseudoadiabat by the adiabatic temperature gradient implied by the H2 equation of state, our results presented here are virtually identical if the binodal pressure is taken to be 2 GPa instead of 4 GPa. This lower pressure for the binodal would imply a younger, less evolved planet.
Hindrance of convection due to the mass load of heavy elements at relatively high temperatures must be included, as described previously for similar circumstances by Misener et al. (2023). By including the Ledoux criterion rather than just the Schwarzschild criterion for convection, the calculations permit development of a radiative layer at the base of the atmosphere that transports the intrinsic heat coming from the magma ocean upward (Guillot, 2010; Misener et al., 2023). The intrinsic luminosity due to the heat emanating from the core to space becomes a determining factor for the structure of the atmosphere. Since radiative diffusion relies on the local luminosity, the temperature gradient due to radiative diffusion includes the radially-dependent luminosity, $L(r)$. At depths well below the radiative-convective boundary, stellar radiation cannot penetrate the atmosphere. Therefore, at these depths the radiative diffusion is given by
@@
We define the radius of the planet using the chord optical depth as formulated by Guillot (2010) coupled with the opacities from Freedman et al. (2014) with metallicities corresponding to the composition of the gas (essentially pure H2 at relevant heights for the outer atmosphere).